OpenStudy (anonymous):

What kind of sequence is the pattern 1, 6, 7, 13, 20, ...? arithmetic sequence exponential sequence geometric sequence recursive sequence

OpenStudy (anonymous):

add the two numbers up to get the next one

Parth (parthkohli):

Do you understand what all sequences actually are?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh that question was probably not for me was it?

OpenStudy (anonymous):

Nope, I don't.

OpenStudy (rajee_sam):

Arihmetic will have a common difference between terms

Parth (parthkohli):

Arithmetic sequence, where \(a_n - a_{n -1} = \rm constant\). Geometric, where \(a_n /a_{n-1} = \rm constant\). (What is an exponential sequence?) Recursive, where \(a_{n + 2} = a_{n+1} + a_{n}\)

terenzreignz (terenzreignz):

<chuckles> recursive sequence, @ParthKohli ?

Parth (parthkohli):

Eh?

terenzreignz (terenzreignz):

\[\large a_{n+2}=a_{n+1}+a_{n}\] Is this a definition or an example? :3

terenzreignz (terenzreignz):

Actually a recursive sequence is one where the terms are defined as functions of their preceding terms... But that's technical stuff :3